OpenStudy (anonymous):

For the parent function y equals the cube root of x what effect does a value of a = -1 have on the graph?

OpenStudy (anonymous):

Does \(y = \sqrt[3]{x -a}\) or \(y = \sqrt[3]{x} +a\) ?

OpenStudy (anonymous):

\[y=\sqrt[3]{x}\]

OpenStudy (anonymous):

what effect does a of -1 have on the graph @KevinOrr

OpenStudy (anonymous):

Well, I don't know where a goes in the function. If it's \(y=\sqrt[3]{x+1}\), then it's `Left 1`. If it's \(y=\sqrt[3]{x}-1\), then it's `Down 1`.